How many moles of pd metal are in each mole of the
Answer #1

Molecular weight of the catalyst = 566.42 gm

At weight of Pd = 106.42 gm

1 mole of catalyst contain 1 mole of Pd

----------------------------------------

0.25 M solution = 0.25 mol in 1000mL solution.

1 mL solution contain 0.00025 moles of Pd

------------------------------------------------------------------------------

moles of 4 -bromo benzoic acid = 0.5/201 = 2.49 *10^-3

moles of phenylboronic acid = 0.37/122 =3.03 *10^-3

the reactant reacts in 1: 1 ratio. As 4-bromo benzoic acid is present in lessed amount (with respect to number of mole), it will be consumed completely and this is the limiting reagent.

ratio of number of moles of Pd: 4-bromobenzoic acid = 0.0025/ 2.49 *10^-3 ~1:1